# Short-video Propagation Influence Rating: A New Real-world Dataset and A New Large Graph Model

### v1

@@ A.4. Statistics of Data
where is the likes in the -th day after the post of the short-video and is the likes increase percentage of the -th day. In Figure 8(a), we plot the likes increase percentage of the -st to -th day after the post. We can observe that the number of likes has become stable since the -th day. The similar trends can be observed for other interactive indicators. These results verify our observation that the interactive information of almost all videos peaks within 2 weeks. Therefore, we can utilize the interactive information later than 2 weeks after the post to annotate the final influence level achieved over a long period.
Length of sample information. We analyze the length of the sample informaiton in this section. We utilize the tokenizer of Qwen2-VL to tokenize the sample information in the json format and compute the number of tokens. In Figure 8(b), we demonstrate the distribution of token lengths in XS-Video. We can find that most samples have 200-500 tokens, while some samples with a large amount of comments have thousands of tokens. Moreover, the distributions of token numbers in the training and test sets are similar. Therefore, learning out-of-distribution patterns of long samples is a challenge in XS-Video.
A.5. Example of Samples
@@
- | - | - | Sum of above edges | 1,667,716,553 |
Appendix B More Details about Method
B.1. Prompt Instruction for Fine-tuning
In supervised language fine-tuning and task-oriented predictor fine-tuning, we combine graph feature and the textual information for SPIR. Assume is the textual information of the video in the json format, the detailed input prompt is as follows:
@@ B.1. Prompt Instruction for Fine-tuning
Appendix C More Details about Experiments
C.1. More Implementation Details
We include more implementation details of our NetGPT in this section. To improve robustness to varying lengths of the input samples while constructing the input instruction, we randomly select 50 comments of a sample if it has more than 50 comments. Moreover, we set the maximum number of tokens as 4000 when tokeninzing the input instructions. In task-oriented predictor fine-tuning, we only train the last 4 layers of the LLM decoder while keeping other decoder layers frozen.
@@ C.1. More Implementation Details
For all LLM and multimodal LLM baselines, we apply the same truncation strategy of inputs in our model. We utilize the LoRA algorithm (Hu et al., 2022) to efficiently optimize these models.

### v2

@@ 3.4 Data Analysis
where is the likes in the -th day after the post of the short-video and is the likes increase percentage of the -th day. In Figure 4(e), we plot the likes increase percentage of the -st to -th day after the post. We can observe that the number of likes has become stable since the -th day. The similar trends can be observed for other interactive indicators. These results verify our observation that the interactive information of almost all videos peaks within 2 weeks. Therefore, we can utilize the interactive information later than 2 weeks after the post to annotate the final influence level achieved over a long period.
Length of sample information. We analyze the length of the sample informaiton in this section. We utilize the tokenizer of Qwen2-VL to tokenize the sample information in the json format and compute the number of tokens. In Figure 4(f), we demonstrate the distribution of token lengths in XS-Video. We can find that most samples have 200-500 tokens, while some samples with a large amount of comments have thousands of tokens. Moreover, the distributions of token numbers in the training and test sets are similar. Therefore, learning out-of-distribution patterns of long samples is a challenge in XS-Video.
3.5 Example of Samples
@@
| - | - | - | Sum of above edges | 1,667,716,553 |
3.6 Short-video Propagation Graph
Utilizing the entities and relations present in XS-Video, we construct a huge propagation graph with the label set . Here is the vertex set, containing 5 million nodes of videos, platforms, topics, titles, descriptions, comments, various interactive indicators, etc. is the edge set, containing 1.7 billion directed edges between nodes, which represent the belonging relations between videos and their attributions, common author relations between videos, common topic relations between videos, chronological relations between samples of the same videos, etc. Specially, is the set of all video nodes and is the set of annotated propagation influence levels of all videos.
Utilizing the entities and relations present in XS-Video, we construct a huge propagation graph consisting of 5.5 million nodes of videos, topics, comments, various interactions, etc., and 1.7 billion directed edges between these nodes. The information about nodes and edges in our graph is shown in Tables III and IV, respectively. The short-video propagation graph is a typical heterogenous graph with both heterogenous nodes and edges. In brief, the nodes include the video nodes, the static attribution nodes of videos, and the dynamic interactive nodes of videos. The edges include the edges from the attribution nodes to videos, the interactive nodes to videos, and the video nodes to video nodes.
4 Method
Short-video propagation network is typical graph-structured data, where edges (such as video-author and video-video) represent complex propagation relations between entries. State-of-the-art methods for analyzing graph-structured data usually adopt Graph Neural Networks (GNNs) [35, 36, 37]. Since Large Language Models (LLMs) have recently shown powerful reasoning ability for textual data, researchers have attempted to adapt LLMs to understand graph-structured data [20, 21, 22]. However, existing research focuses on textual graphs where the information of nodes and edges can be transformed into text, thereby directly leveraging the textual reasoning ability of LLMs. These methods cannot be directly applied to our heterogeneous propagation graph of short-videos. To analyze the short-video propagation graph, where nodes can be texts, times, scalars, and videos, we propose a novel Large Graph Model (LGM) named NetGPT NetGPT fuses open-source LLM (i.e., Qwen2-VL [18] in our experiments) and heterogeneous GNN to effectively comprehend and analyze the short-video propagation graph. Figure 6 illustrates the framework of our three-stage method.
4.1 Heterogeneous Graph Pretraining
Given the short-video propagation graph formulated in Section 3.6 and Tables III-IV, we first extract the raw features of all heterogenous nodes in . For each node of the video modality, we utilize the ViT [38] in Qwen2-VL-7B-Instruct444https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct to extract the video feature with an average pooling operation, as follows:
where is the average pooling operation that computes the average feature of the spatial-temporal features outputted by the ViT. For each node of the text modality (including platforms, topics, titles, and descriptions), we utilize the RoBERTa [39, 40] of the Chinese-RoBERTa-wwm-ext-large version555https://huggingface.co/hfl/chinese-roberta-wwm-ext-large to extract the text feature, as follows:
For each node of the time modality (including times and ctimes), as is formulated in the timestamp format, we utilize Sinusoidal Encoding [41] to compute the time feature, as follows:
where is the output feature of the node , is the dimension of the feature, and is a 2-layer RGCN [42]. Then, we add a 1-layer predictor layer and pretrain the GNN with the predictor by Smooth L1 loss [43], as follows:
4.2 Supervised Language Fine-tuning
where and are trainable parameters, and is the token embedding dimension of the LLM. Next, we construct the input instruction for LLM, which is “You are a helpful assistant. Graph: . Please read the following json format information about the short-video and predict the final propagation influence level (levels 0 9) of the short video: ”, where is the textual information of the sample (as shown in Figure 2 except the influence level) in the json format. More details of our instruction are introduced in Appendix 4.3. Then, we obtain the input token embeddings of , where the embedding of is computed by Equations 8 and 10, and the embeddings of other textual tokens are obtained by the embedding layer of the LLM. The ground truth output of this stage is constructed as response =“The influence level of this short-video is ”, where is the ground truth influence level of . The optimization objective of this stage is maximizing the generation probability of with the input , as follows:
where is the generation probability function of the LLM. Following the original implementation of Qwen2-VL [18], we leverage the cross-entropy loss to optimize the Equation 11. Specially, to preserve the pretrained knowledge and avoid the model corruption introduced by fusing GNN and LLM, we only train the parameters and of the graph feature projector, while keeping all parameters of the pretrained GNN and LLM frozen.
4.3 Prompt Instruction for Fine-tuning
In supervised language fine-tuning and the following task-oriented predictor fine-tuning, we combine the graph feature and the textual information for SPIR. Assume is the textual information of the video in the json format, the detailed input prompt is as follows:
@@ 4.3 Prompt Instruction for Fine-tuning
4.4 Task-oriented Predictor Fine-tuning
5 Experiments
5.1 Experimental Setup
Baseline methods. We adopt 10 state-of-the-art baseline methods, including 4 GNNs, 4 LLMs, and 2 multimodal LLMs. The adopted 4 GNN baselines are GCN [44], HAN [45], HetSANN [35], and RGCN [46]. The adopted 4 LLM baselines are Mistral-7B-Instruct-v0.3 [47], InternLM2.5-7B [48], Llama-3.1-8B-Instruct [49], and Qwen2.5-7B-Instruct [50]. We further adopt 2 multimodal LLM baselines that can comprehend the video modality, i.e., Llava-Next-Video-7B [19] and Qwen2-VL-7B-Instruct [18]. To conduct a fair comparison, we use the same input graph, predictor module, and optimization loss introduced in Section 4.1 for all GNN baselines. For LLM and multimodal LLM baselines, we use the same input instructions, output response, and optimization loss introduced in Section 4.2, where the information of the short-video is organized in the json format. Additionally, we input the video data into multimodal LLM baselines.
Implementation details. We adopt a 2-layer RGCN as the graph encoder and Qwen2-VL-7B-Instruct as the LLM backbone in our model. We utilize the DeepSpeed [51] framework and ZeRO-2 optimizer [52] to train our model on 8 Nvidia A800 80G GPUs. is set to 1024. To improve robustness to varying lengths of the input samples while constructing the input instruction, we randomly select 50 comments of a sample if it has more than 50 comments. Moreover, we set the maximum number of tokens as 4000 when tokeninzing the input instructions. In task-oriented predictor fine-tuning, we only train the last 4 layers of the LLM decoder while keeping other decoder layers frozen.
@@ 5.1 Experimental Setup
For all LLM and multimodal LLM baselines, we apply the same truncation strategy of inputs in our model. We utilize the LoRA algorithm [53] to efficiently optimize these models.
5.2 Metrics
Accuracy (ACC). Following widely-used evaluation for the classification task [54, 55, 56], we compute the Accuracy (ACC) of the predicted results as follows:
where is the rounding function.
Mean Squared Error (MSE). Following widely-used evaluation for the regression task [57, 55, 1], we compute the Mean Squared Error (MSE) of the predicted results as follows:
| Model | Input | ACC | MSE | MAE |
| GCN | Graph | 0.4474 | 1.0623 | 0.7599 |
| HAN | Graph | 0.2619 | 2.6666 | 1.2724 |
| HetSANN | Graph | 0.5078 | 0.8917 | 0.6803 |
| RGCN | Graph | 0.6313 | 0.7801 | 0.5844 |
| Mistral-v0.3 | Text | 0.5387 | 2.1000 | 0.8123 |
| InternLM2.5 | Text | 0.5268 | 2.1110 | 0.8064 |
| Llama-3.1 | Text | 0.5290 | 2.1215 | 0.8081 |
| Qwen2.5 | Text | 0.5469 | 2.0820 | 0.7688 |
| Llava-Next-Video | Text+Video | 0.5694 | 1.8503 | 0.7315 |
| Qwen2-VL | Text+Video | 0.5884 | 1.6820 | 0.6629 |
| NetGPT (Ours) | Graph+Text | 0.6777 | 0.7169 | 0.5457 |
5.3 Results and Discussions
The experimental results on our XS-Video dataset are shown in Table V. From the results, we have the following observations:
| Model | ACC | MSE | MAE |
| NetGPT-V | 0.6564 | 0.7415 | 0.5669 |
| NetGPT-VV | 0.5889 | 0.8935 | 0.6386 |
| NetGPT-IV | 0.2805 | 2.2839 | 1.1688 |
| NetGPT-CV | 0.6464 | 0.8249 | 0.5868 |
| NetGPT-SLF | 0.6552 | 0.7366 | 0.5605 |
| NetGPT-2B | 0.6695 | 0.7307 | 0.5563 |
| NetGPT (Ours) | 0.6777 | 0.7169 | 0.5457 |
(a) Long-term prediction
(b) Median-term prediction
(c) Short-term prediction
5.4 Ablation Study
As shown in Table VI, we have the following observations: (1) NetGPT-V performs worse than NetGPT, which shows the importance of considering the video content information. (2) NetGPT-VV, NetGPT-IV, and NetGPT-CV perform worse than NetGPT, which shows the edges from video nodes, interactive information nodes, and comment content nodes to video nodes are significant. (3) NetGPT-SLF underperforms NetGPT, which verifies the effectiveness of our three-stage training mechanism. (4) With a smaller LLM backbone, NetGPT-2B performs worse than NetGPT, which indicates the effectiveness of leveraging pretrained knowledge of larger LLMs.
5.5 Results of Short- and Long-term Prediction
(a) ACC in training procedure
(b) MSE in training procedure
5.6 Convergence on Different Platforms
6 Discussion and Conclusion